# A logic of co-valuations

### v1

@@ 5.3. The Polish space $\rm{Mod}(L)$ of infinite compact $\omega$-structures in a countable signature $L$
The proof of Theorem 6.6 gives that Fraïssé limits are $\omega$-categorical relative to the class of structures that can be obtained from the respective category. For example, the atomic model of the pseudo-arc is the unique chainable $\omega$-compact model of its theory.
As a matter of fact, we can explicitely characterize $\omega$-structures homeomorphic with the arc that are Fraïssé limits of $\mathbf{I}$. Recall that a taut cover of a topological space $X$ is a cover $C$ such that ${\rm{cl}(c)}\cap{\rm{cl}(c^{\prime})}\neq\emptyset$ implies $c\cap c^{\prime}\neq\emptyset$, for $c,c^{\prime}\in C$. We will say that a co-valuation $v$ is connected if every $x\in[v]$ is connected. In particular, connected open co-valuations on the arc are built out of open intervals.
@@
We consider only the former case; the other one is analogous. Since $C$ fragments $v_{i}]$, there must be $y^{\prime}=(a_{y^{\prime}},b_{x})\in C$ with

### v2

@@ 5.3. The Polish space $\rm{Mod}(L)$ of infinite compact $\omega$-structures in a countable signature $L$
The proof of Theorem 6.6 gives that Fraïssé limits are $\omega$-categorical relative to the class of structures that can be obtained from the respective category. For example, the atomic model of the pseudo-arc is the unique chainable $\omega$-compact model of its theory. Also, Proposition 4.27 says that the ordered arc is the unique metrizable $\omega$-compact model of its theory.
As a matter of fact, we can explicitely characterize $\omega$-structures homeomorphic with the arc that are Fraïssé limits of $\mathbf{I}$. Recall that a taut cover of a topological space $X$ is a cover $C$ such that ${\rm{cl}(c)}\cap{\rm{cl}(c^{\prime})}\neq\emptyset$ implies $c\cap c^{\prime}\neq\emptyset$, for $c,c^{\prime}\in C$. We will say that a co-valuation $v$ is connected if every $x\in[v]$ is connected. In particular, connected open co-valuations on the arc are consist of open intervals.
@@
We consider only the former case, the other one is analogous. Since $C$ fragments $v_{i}]$, there must be $y^{\prime}=(a_{y^{\prime}},b_{x})\in C$ with